OpenStudy (anonymous):

At a specialty nut store, almonds cost $4.30 per pound and walnuts are priced at $3.75 per pound. How many pounds of walnuts are in a 12-pound mixture priced at $47.75? check my answer?

OpenStudy (anonymous):

answer choices are: 4 5 6 7 I got 5

OpenStudy (amistre64):

lets see if that fits

OpenStudy (amistre64):

5lbs + 7lbs = 12lbs 3.75(5) + 4.30(7) =? 47.75 ; no

OpenStudy (anonymous):

The correct answer is 7.

OpenStudy (anonymous):

thank you, @escrite, but i would like to know HOW to get that answer

OpenStudy (amistre64):

3.75(5) + 4.30(7) is a close guess, that we can adjust if need be 3.75(5) + 4.30(7) = 48.85 ... its too heavy 3.75(6) + 4.30(6) = 48.30 ... its too heavy 3.75(7) + 4.30(5) = 47.75 just right i think you did it correctly but got the a and ws mixed up

OpenStudy (anonymous):

oooohhhhh.... yup, that was my mistake. thanks!

OpenStudy (amistre64):

substitution for this is simple enough w + a = 12 : a = 12-w 3.75w + 4.30a = 47.75 3.75w + 4.30(12-w) = 47.75 solve for w